# Solving stochastic climate-economy models: A deep least-squares Monte Carlo approach

### v1

@@ 5 Conclusions
This paper presents an advanced approach to modeling recursive stochastic climate-economy models using a deep least-squares Monte Carlo (LSMC) method. The method’s flexibility allows for the application to various types of uncertainties, including parametric and stochastic process uncertainties. The integration of deep neural networks enables the handling of high-dimensional models in a tractable manner and within a reasonable computational budget, thus making stochastic climate-economy models more accessible to researchers and policymakers. The methodology and findings presented here provide a solid foundation for future work in this vital area of research.
Future research should explore the incorporation of Bayesian learning mechanisms to update probabilities as more information becomes available over time. Since our approach can manage high-dimensional stochastic shocks, a natural next step is to study the impact of multi-dimensional probability distributions whose marginals are correlated. Additionally, we aim to apply our method to the study of climate tipping points as well as the Regional Integrated model of Climate Change and the Economy (RICE) of Nordhaus and Yang (1996). These future steps could further refine the model’s predictions and enhance its policy relevance.
It is important to note that IAMs, and the DICE model in particular, have limitations in the model structure and model parameters which are debated in the literature, see e.g. discussions in Pindyck (2017). The incorporation of uncertainties into these models is an important improvement. Our approach demonstrates significant advancements in modeling and solving complex stochastic climate-economy models. By capturing a wide range of uncertainties and leveraging advanced computational techniques, we contribute to the development of more robust and reliable tools for climate policy analysis. The continued evolution of these models will be critical in supporting effective and sustainable climate action in the years to come, and the deep least-squares Monte Carlo method provides a useful tool to solve stochastic climate-economy models.
@@ Acknowledgements
Aleksandar Arandjelović acknowledges support from the International Cotutelle Macquarie University Research Excellence Scholarship, as well as from the Institute of Statistical Mathematics, Japan. A significant part of work on this project was carried out while Aleksandar Arandjelović was affiliated with the research unit Financial and Actuarial Mathematics (FAM) at TU Wien in Vienna, Austria as well as the Department of Actuarial Studies and Business Analytics at Macquarie University in Sydney, Australia. Pavel Shevchenko and Tor Myrvoll acknowledge travel support from the Institute of Statistical Mathematics, Japan. This work was presented at the international workshop “Stochastic Modelling in Climate Risk” in November 2023 (ISM Japan), the international workshop “Advances in Risk Modelling and Applications to Finance and Climate Risk” in July 2024 (WU Wien, Austria), and at invited seminars at the University of New South Wales, the Australian National University, Vienna University of Technology and Mahidol University in 2024. The authors are grateful for the constructive comments received from participants of these events.
Literatur
  • Ackerman et al. (2009) Frank Ackerman, Stephen J. DeCanio, Richard B. Howarth, and Kristen Sheeran. Limitations of integrated assessment models of climate change. Climatic Change, 95(3):297–315, 2009.
  • Aïd et al. (2014) René Aïd, Luciano Campi, Nicolas Langrené, and Huyên Pham. A probabilistic numerical method for optimal multiple switching problems in high dimension. SIAM Journal on Financial Mathematics, 5(1):191–231, 2014.
  • Andréasson and Shevchenko (2022) Johan G. Andréasson and Pavel V. Shevchenko. A bias-corrected least-squares Monte Carlo for solving multi-period utility models. European Actuarial Journal, 12(1):349–379, 2022.
  • Andréasson and Shevchenko (2024) Johan G. Andréasson and Pavel V. Shevchenko. Optimal annuitisation, housing and reverse mortgage in retirement in the presence of a means-tested public pension. European Actuarial Journal, 2024. https://doi.org/10.1007/s13385-024-00379-3.
  • Belomestny (2011) Denis Belomestny. Pricing Bermudan options by nonparametric regression: Optimal rates of convergence for lower estimates. Finance and Stochastics, 15:655–683, 2011.
  • Belomestny et al. (2010) Denis Belomestny, Anastasia Kolodko, and John Schoenmakers. Regression methods for stochastic control problems and their convergence analysis. SIAM Journal on Control and Optimization, 48(5):3562–3588, 2010.
  • Cai and Lontzek (2019) Yongyang Cai and Thomas S. Lontzek. The social cost of carbon with economic and climate risks. Journal of Political Economy, 127(6):2684–2734, 2019.
  • Cybenko (1989) George Cybenko. Approximation by superpositions of a sigmoidal function. Mathematics of Control, Signals and Systems, 2(4):303–314, 1989.
  • Dvoretzky (1956) Aryeh Dvoretzky. On stochastic approximation. In Proceedings of the Third Berkeley Symposium on Mathematical Statistics and Probability, 1954–1955, vol. I, pages 39–55. University of California Press, Berkeley and Los Angeles, Calif., 1956.
  • Fleming and Soner (2006) W.H. Fleming and H.M. Soner. Controlled Markov Processes and Viscosity Solutions, volume 25 of Stochastic Modelling and Applied Probability. Springer, New York, second edition, 2006.
  • Friedl et al. (2023) Aleksandra Friedl, Felix Kübler, Simon Scheidegger, and Takafumi Usui. Deep uncertainty quantification: With an application to integrated assessment models. Technical report, Working Paper University of Lausanne, 2023.
  • Gillingham et al. (2015) Kenneth Gillingham, William D. Nordhaus, David Anthoff, Geoffrey Blanford, Valentina Bosetti, Peter Christensen, Haewon McJeon, John Reilly, and Paul Sztorc. Modeling uncertainty in climate change: A multi-model comparison. Technical report, National Bureau of Economic Research, 2015.
  • Grubb et al. (2021) Michael Grubb, Claudia Wieners, and Pu Yang. Modeling myths: On DICE and dynamic realism in integrated assessment models of climate change mitigation. Wiley Interdisciplinary Reviews: Climate Change, 12(3):e698, 2021.
  • Hornik (1991) Kurt Hornik. Approximation capabilities of multilayer feedforward networks. Neural Networks, 4(2):251–257, 1991.
  • Ikefuji et al. (2020) Masako Ikefuji, Roger J. A. Laeven, Jan R. Magnus, and Chris Muris. Expected utility and catastrophic risk in a stochastic economy-climate model. Journal of Econometrics, 214(1):110–129, 2020.
  • Interagency Working Group on Social Cost of Greenhouse Gases (2016) Interagency Working Group on Social Cost of Greenhouse Gases. Technical support document: Social cost of carbon for regulatory impact analysis under executive order 12866. Technical report, United States Government, 2016.
  • Jensen and Traeger (2014) Svenn Jensen and Christian P. Traeger. Optimal climate change mitigation under long-term growth uncertainty: Stochastic integrated assessment and analytic findings. European Economic Review, 69:104–125, 2014.
  • Kelly and Kolstad (1999) David L. Kelly and Charles D. Kolstad. Bayesian learning, growth, and pollution. Journal of Economic Dynamics and Control, 23(4):491–518, 1999.
  • Kelly and Kolstad (2001) David L. Kelly and Charles D. Kolstad. Solving infinite horizon growth models with an environmental sector. Computational Economics, 18:217–231, 2001.
  • Kharroubi et al. (2014) Idris Kharroubi, Nicolas Langrené, and Huyên Pham. A numerical algorithm for fully nonlinear HJB equations: An approach by control randomization. Monte Carlo Methods and Applications, 20(2):145–165, 2014.
  • Kiefer and Wolfowitz (1952) Jack Kiefer and Jacob Wolfowitz. Stochastic estimation of the maximum of a regression function. Annals of Mathematical Statistics, 23(3):462–466, 1952.
  • Kratsios (2021) Anastasis Kratsios. The universal approximation property: Characterization, construction, representation, and existence. Annals of Mathematics and Artificial Intelligence, 89(5):435–469, 2021.
  • Leach (2007) Andrew J. Leach. The climate change learning curve. Journal of Economic Dynamics and Control, 31(5):1728–1752, 2007.
  • Longstaff and Schwartz (2001) Francis A. Longstaff and Eduardo S. Schwartz. Valuing American options by simulation: A simple least-squares approach. The Review of Financial Studies, 14(1):113–147, 2001.
  • Lontzek et al. (2015) Thomas S. Lontzek, Yongyang Cai, Kenneth L. Judd, and Timothy M. Lenton. Stochastic integrated assessment of climate tipping points indicates the need for strict climate policy. Nature Climate Change, 5(5):441–444, 2015.
  • Nordhaus (2018) William D. Nordhaus. Projections and uncertainties about climate change in an era of minimal climate policies. American Economic Journal: Economic Policy, 10(3):333–360, 2018.
  • Nordhaus and Yang (1996) William D. Nordhaus and Zili Yang. A regional dynamic general-equilibrium model of alternative climate-change strategies. The American Economic Review, 86(4):741–765, 1996.
  • Nordhaus et al. (1992) William D. Nordhaus et al. The ‘DICE’ model: background and structure of a dynamic integrated climate-economy model of the economics of global warming. Technical report, Cowles Foundation for Research in Economics, Yale University, 1992.
  • Pindyck (2017) Robert S. Pindyck. The use and misuse of models for climate policy. Review of Environmental Economics and Policy, 11:100–114, 2017.
  • Robbins and Monro (1951) Herbert Robbins and Sutton Monro. A stochastic approximation method. Annals of Mathematical Statistics, 22(3):400–407, 1951.
  • Rumelhart et al. (1986) D. E. Rumelhart, G. E. Hinton, and R. J. Williams. Learning representations by back-propagating errors. Nature, 323(6088):533–536, 1986.
  • Shevchenko et al. (2022) Pavel V. Shevchenko, Daisuke Murakami, Tomoko Matsui, and Tor A. Myrvoll. Impact of COVID-19 type events on the economy and climate under the stochastic DICE model. Environmental Economics and Policy Studies, 24:459–476, 2022.
  • Sobol’ (1967) Ilya M. Sobol’. On the distribution of points in a cube and the approximate evaluation of integrals. USSR Computational Mathematics and Mathematical Physics, 7(4):86–112, 1967.
  • Sobol’ (2001) Ilya M. Sobol’. Global sensitivity indices for nonlinear mathematical models and their Monte Carlo estimates. Mathematics and Computers in Simulation, 55(1–3):271–280, 2001.
  • Traeger (2014) Christian P. Traeger. A 4-stated DICE: Quantitatively addressing uncertainty effects in climate change. Environmental and Resource Economics, 59(1):1–37, 2014.
  • Tsitsiklis and Van Roy (2001) John N. Tsitsiklis and Benjamin Van Roy. Regression methods for pricing complex American-style options. IEEE Transactions on Neural Networks, 12(4):694–703, 2001.
  • Weitzman (2011) Martin L. Weitzman. Fat-tailed uncertainty in the economics of catastrophic climate change. Review of Environmental Economics and Policy, 5(2):275–292, 2011.

### v2

@@ 5 Conclusions
This paper presents an advanced approach to modeling recursive stochastic climate-economy models using a deep least-squares Monte Carlo (LSMC) method. The method’s flexibility allows for the application to various types of uncertainties, including parametric and stochastic process uncertainties. The integration of deep neural networks enables the handling of high-dimensional models in a tractable manner and within a reasonable computational budget, thus making stochastic climate-economy models more accessible to researchers and policymakers. The methodology and findings presented here provide a solid foundation for future work in this vital area of research. More broadly, the resulting scenario distributions can serve as inputs for downstream climate-finance and actuarial applications; see, e.g., arandjelovicShevchenko2025netzero; arandjelovicShevchenko2025reserves.
Future research should explore the incorporation of Bayesian learning mechanisms to update probabilities as more information becomes available over time. Since our approach can manage high-dimensional stochastic shocks, a natural next step is to study the impact of multi-dimensional probability distributions whose marginals are correlated. Another natural extension is to incorporate risk-sensitive recursive preferences that separate IES and risk aversion (e.g. Epstein–Zin), as in cai2019, to quantify how preference specification alters precautionary motives and the SCC under the same uncertainty blocks. Additionally, we aim to apply our method to the study of climate tipping points as well as the Regional Integrated model of Climate Change and the Economy (RICE) of nordhaus1996. These future steps could further refine the model’s predictions and enhance its policy relevance.
Extending from the global DICE model to the multi-region RICE model (nordhaus1996) increases the dimension of the economic state (e.g. regional capital stocks) and the number of controls (regional consumption and mitigation). In the deep LSMC algorithm, however, the regression step targets the continuation-value conditional expectation as a function of the post-decision state; hence the effective regression input dimension scales with the number of endogenous state variables rather than with the number of controls. The dominant costs are (i) simulating training samples and (ii) fitting the regression model at each time step, both of which are parallel and can be distributed across CPU/GPU resources. While larger training sets will be required as the effective dimension increases (and regional structure should be exploited), preliminary experiments on a 12-region prototype under a single stochastic shock indicate that the approach remains numerically stable; a full multi-shock implementation is left for future work.
It is important to note that IAMs, and the DICE model in particular, have limitations in the model structure and model parameters which are debated in the literature, see e.g. discussions in pindyck2017use. The incorporation of uncertainties into these models is an important improvement. Our approach demonstrates significant advancements in modeling and solving complex stochastic climate-economy models. By capturing a wide range of uncertainties and leveraging advanced computational techniques, we contribute to the development of more robust and reliable tools for climate policy analysis. The continued evolution of these models will be critical in supporting effective and sustainable climate action in the years to come, and the deep least-squares Monte Carlo method provides a useful tool to solve stochastic climate-economy models.
@@ Acknowledgements
Aleksandar Arandjelović acknowledges support from the International Cotutelle Macquarie University Research Excellence Scholarship, as well as from the Institute of Statistical Mathematics, Japan. A significant part of work on this project was carried out while Aleksandar Arandjelović was affiliated with the research unit Financial and Actuarial Mathematics (FAM) at TU Wien in Vienna, Austria as well as the Department of Actuarial Studies and Business Analytics at Macquarie University in Sydney, Australia. Pavel Shevchenko and Tor Myrvoll acknowledge travel support from the Institute of Statistical Mathematics, Japan. This work was presented at the international workshop “Climate Finance & Risk” in November 2024 (ISM Japan), the international workshop “Advances in Risk Modelling and Applications to Finance and Climate Risk” in July 2024 (WU Wien, Austria), “Quantitative Methods in Finance” international conference in December 2025 (Sydney, Australia), and at invited seminars at the University of New South Wales, the Australian National University, Vienna University of Technology and Mahidol University in 2024. The authors are grateful for the constructive comments received from participants of these events.